OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions: Through (-2,11); perpendicular to the line 4x-8y=1 Detailed steps if necessary, please and many thanks.

OpenStudy (anonymous):

I think you can use <4,-8> as a vector... then x=-2+4t and y = 11-8t

OpenStudy (anonymous):

do you have the answer and need to understand better? It always helps to graph on paper...

OpenStudy (amistre64):

just swap coeffs and switch the operator

OpenStudy (amistre64):

..and readjust for a new constant 4x - 8y = 1 8x + 4y = C -2 11 -16 + 44 =28 8x + 4y = 28

OpenStudy (anonymous):

What level math is this? I feel like I am out of my league. I do not understand the terminology.